The second side of a triangle is four more than the first side, and the third side is five more than the second. If the product of the first and third sides is added to the product of the second and thirds sides, the result is 124 more than twice the square of the first side. find the perimeter of the triangle.
Answer by VFBundy(438)

1st side = x
2nd side = x + 4
3rd side = (x + 4) + 5 = x + 9

x(x + 9) + (x + 4)(x + 9) = 2x² + 124

x² + 9x + x² + 13x + 36 = 2x² + 124

2x² + 22x + 36 = 2x² + 124

22x + 36 = 124

22x = 88

x = 4

1st side = x = 4
2nd side = x + 4 = 8
3rd side = x + 9 = 13

The perimeter of the triangle is 25.